# Alberta Wells Dataset: Pinpointing Oil and Gas Wells from Satellite Imagery

### v1

@@ 4.2 Object Detection
All object detection models are trained using a ResNet50 backbone, except for SSD Lite, which is trained with a MobileNet backbone. The batch size is set to 256 for Faster R-CNN and FCOS and 512 for RetinaNet and SSD Lite. We used a cosine annealing scheduler [40] and trained all models for 120 epochs. All models are optimized using the AdamW optimizer [44].
@@
4.3 Results & Analysis
@@
Architecture Backbone IoU F1 Score Precision Recall
U-Net ResNet50 58$\pm$0.5 61.9$\pm$0.8 90.2$\pm$2.2 62.3$\pm$1.6
ResNext50 58.2$\pm$0.2 62.1$\pm$0.3 88.2$\pm$3.5 63.6$\pm$1.7
SE_ResNet50 58.9$\pm$0.7 62.9$\pm$0.7 88.8$\pm$1.6 64.4$\pm$1.4
EfficientNetB6 60.4$\pm$0.3 64.8$\pm$0.4 87.8$\pm$0.4 66.3$\pm$0.3
PAN ResNet50 57.8$\pm$0.8 61.5$\pm$0.9 89.3$\pm$1.22 61.5$\pm$0.93
DeepLabV3+ 56.8$\pm$0.7 60.6$\pm$0.7 89.4$\pm$1.3 61.8$\pm$1.1
Segformer mit-b0-ade 57.6$\pm$0.5 61.3$\pm$0.6 82.6$\pm$2.9 69.2$\pm$2.1
UperNet swin small 59.9$\pm$0.7 64.2$\pm$0.7 80.6$\pm$0.5 73.1$\pm$0.1
Our tasks involve identifying a roughly circular well region with a 90m diameter in real life, which translates to less than 30 pixels in satellite imagery due to resizing and other augmentations. This poses a challenge for machine learning models given the heterogeneous nature of the background, including various similarly shaped and sized features of the natural and built environment. Additionally, vegetation can occlude wells in RGB channels, highlighting the importance of near-infrared imagery for guiding the model. The wells themselves also vary somewhat in shape and can be in various states of disrepair as a result of differing ages and maintenance.
4.3.1 Binary Segmentation
For the binary segmentation task framing, we train both CNN-based and Transformer-based backbones, considering the prevalent imbalance in the image data due to the small size of wells.
Among our models, as shown in Table 4, the traditional U-Net with EfficientNetB6 backbone performs the best, with CNN-based models showing the highest IOU of $60.4\pm 0.3$ and F1-Score of $64.8\pm 0.4$. While a ResNet50-based backbone achieves the highest Precision of $90.2\pm 2.2$, indicating more accurate predictions of well instances compared to other models. Precision, which reflects the accuracy of our positive detections compared to the ground truth, is crucial. However, a high recall value ensures the model captures most actual well instances, reducing the risk of missing important information. Thus, the Uper-Net model with the highest recall value of $73.1\pm 0.1$, which excels at capturing global context information, appears a good candidate for this task given a decent precision score. However, taking into account both precision and recall, U-Net with EfficientNetB6 backbone perform well, suggesting the utility of a larger backbone with a bigger receptive field.
@@ 4.3.1 Binary Segmentation
Architecture Backbone IoU0.1 IoU0.3 IoU0.5 mAP50 mAP50:95
RetinaNet ResNet50 24.58$\pm$0.11 43.07$\pm$0.8 59.79$\pm$0.36 0.18$\pm$0.28 0.72$\pm$1.12
FasterRCNN 36.79$\pm$1.07 46.95$\pm$0.66 61.29$\pm$0.35 5.2$\pm$1.0 19.12$\pm$3.41
FCOS 34.79$\pm$0.99 48.51$\pm$0.59 62.66$\pm$0.43 9.67$\pm$1.47 30.46$\pm$3.11
SSD Lite MobileNet 33.91$\pm$0.18 50.30$\pm$0.08 65.07$\pm$0.03 9.76$\pm$0.39 25.14$\pm$0.66
4.3.2 Binary Object Detection
@@
The observed gap in performance is likely due to the small size of the wells. It is well-known that single-stage CNN architectures (such as FCOS and SSD) often demonstrate better performance on small object detection than two-stage methods (such as Faster R-CNN), and this aligns with our observations. The exception here is the single-stage method RetinaNet, which, although it has comparable IoU scores, struggles to detect wells accurately. SSD Lite stands out with the highest IoU0.5 score of $65.07\pm 0.03$ and IoU0.3 score of $50.3\pm 0.08$. Whereas all models are quite similar in terms of IoU0.1, the highest score by FasterRCNN is $36.79\pm 1.07$. Thus, SSD Lite and FCOS excel in localization, especially at higher IoU thresholds, while Faster R-CNN is adept at detecting objects with minimal overlap. All models demonstrate low performance in terms of mAP50, which assesses precision-recall trade-off and detection accuracy at an IoU threshold of 0.5. FCOS achieves the of $9.67\pm 1.47$ while SSD Lite achieves a score of $9.76\pm 0.39$. This may be due to these models not producing region proposals confidently enough, especially in instances with a large number of wells. Whereas over a broader evaluation with mAP50:95 which averages precision across IoU thresholds from 0.5 to 0.95. All models apart from RetinaNet provide much better results, with FCOS achieving the highest score of $30.46\pm 3.11$, indicating a decent performance in the identification of well instances.
@@ 6 Conclusion
In this paper, we present the first large-scale benchmark dataset aimed at identifying oil and gas wells, with a focus on abandoned and suspended wells, which are a significant source of greenhouse gases and other pollutants. We combine high-resolution imagery, an extensive database of well locations, and expert verification to create the Alberta Wells Dataset. We frame well identification both in terms of object detection and binary segmentation and evaluate the performance of a wide range of popular deep learning methods on these tasks. We find that the UNet model (with a EfficientNetB6 backbone), in particular, represents the most promising baseline for the binary segmentation task, while for object detection, all models demonstrate more mixed results, with Single Stage Models (such as FCOS and SSD) providing a relatively promising baseline. These results show that the Alberta Wells Dataset represents both a challenging as well as a societally impactful set of tasks.
The value added by the dataset is twofold. First, most global fossil fuel-producing regions do not have databases of well locations comparable to that provided by AER. Alberta’s varied landscape provides an ideal setting for developing algorithms for the detection of wells, which can then be used directly in other locations or fine-tuned. Second, even in Alberta, the list developed by AER is not comprehensive, and many abandoned wells are believed to be missing. Our work can provide candidate locations for domain experts to examine so as to determine the true number and locations of abandoned wells in Alberta.
We hope that our work may be of use to policymakers and other stakeholders involved in climate action and environmental protection according to the following envisioned steps:
  • Use the Alberta Wells Dataset to train algorithms for pinpointing well locations.
  • Run these algorithms at scale across a broader region of interest, comparing against any existing databases to identify those wells that may be undocumented.
  • Flag abandoned wells for plugging, prioritizing those identified as super-emitters.
We believe that the scalability of machine learning tools for remote sensing will make them an invaluable tool in pinpointing and mitigating the global environmental impact of abandoned oil and gas wells.
@@ 7 Acknowledgments and Disclosure of Funding
This research was made possible via Planet Labs’ academic subscription program. We are grateful to Planet Labs and Tim Elrick for input in using this data. Funding was provided by a McGill Tomlinson-Wares grant, Mila - Quebec AI Institute, the Québec Ministère de l’Économie, de l’Innovation et de l’Énergie, and the Canada CIFAR AI Chairs program. The computing resources were provided by Mila - Quebec AI Institute, the Digital Research Alliance of Canada, and NVIDIA Corporation. Special thanks to Francis Pelletier for helpful feedback on our data collection pipeline and code.
Appendix A Dataset Information
@@
Dataset
Split
No of
Samples
No of Wells
in Split
Original HDF5
File Size (in Gb)
Compressed .tar.gz
File Size (in Gb)
Train 167436 194231 322 100
Validation 9463 8243 19 5.7
Test 11789 10973 24 7.1
Total 188688 213447 365 112.8
A.1 Dataset Structure
We provide training, validation, and testing sets, split using our proposed algorithm (as described in Section 3.2 of the main paper) to create a well-distributed dataset.
The proposed method aims to create smaller regions of well concentration by clustering the centroids of patches. These regions are designed to be (a) mutually non-intersecting, (b) part of a larger geographic region by clustering the centroids of the initial clusters, and (c) containing a similar distribution of non-well patches within the same region. This approach ensures that the training, validation, and test sets include representations from all geographic regions, providing a diverse and comprehensive evaluation. Thus, the dataset represents various geographical regions and offers a diverse benchmark for evaluation and testing.
@@ A.1 Dataset Structure
A.2 Dataset File Directory Structure
@@
    |---description: Alberta Wells Dataset:
                     Pinpointing Oil and Gas Wells
@@
A.3 Structure of Dataset Directory
@@
    |---description: Alberta Wells Dataset:
                     Pinpointing Oil and Gas Wells
@@
A.4 Dataset Size & Distribution of Samples
The proposed dataset comprises over 94,000 patches of satellite imagery containing wells, with a total of 188,000 patches sourced from Planet Labs [16]. This dataset covers more than 213,000 individual wells. Details about the distribution of the number of patches, wells present, and dataset split sizes are provided in Table 6, with the distribution of the number of wells per sample being described in Table 7. We also include an equal number of images that contain no wells in each dataset split.

### v2

@@ 4.2 Object Detection
All object detection models are trained using a ResNet50 backbone, except for SSD Lite, which is trained with a MobileNet backbone. The batch size is set to 256 for Faster R-CNN and FCOS and 512 for RetinaNet and SSD Lite. We used a cosine annealing scheduler [40] and trained all models for 120 epochs. All models are optimized using the AdamW optimizer.
@@
Figure 7: A sample image patch from our dataset includes examples with no wells, two wells, and multiple wells. Additionally, we present qualitative results with predictions generated by our Segmentation U-Net (EfficientNet-B6) and Detection FCOS model.
5 Results
Our tasks involve identifying a roughly circular well region with a 90m diameter in real life, which translates to less than 30 pixels in satellite imagery due to resizing and other augmentations.
This poses a challenge for machine learning models given the heterogeneous nature of the background, including various similarly shaped and sized features of the natural and built environment. Additionally, vegetation can occlude wells in RGB channels, highlighting the importance of near-infrared imagery for guidance. The wells themselves also vary somewhat in shape and can be in various states of disrepair as a result of differing ages and maintenance.
5.1 Binary Segmentation
For the binary segmentation task framing, we train Models (from scratch) using both CNN-based and Transformer-based backbones, considering the prevalent imbalance in the image data due to the small size of wells. Although we did use 3-dimensional, ImageNet initialized weights of the backbone but modified the initial layers afterwards to support 4-dimensional multispectral images.
Among our models, as shown in Table 4, the traditional U-Net with EfficientNetB6 backbone performs the best, with CNN-based models showing the highest IOU of $60.4\pm 0.3$ and F1-Score of $64.8\pm 0.4$. While a ResNet50-based backbone achieves the highest Precision of $90.2\pm 2.2$, indicating more accurate predictions of well instances compared to other models. Precision, which reflects the accuracy of our positive detections compared to the ground truth, is crucial. However, a high recall value ensures the model captures most actual well instances, reducing the risk of missing important information. Thus, the Uper-Net model with the highest recall value of $73.1\pm 0.1$, which excels at capturing global context information, appears a good candidate for this task, given a decent precision score.
Taking into account both precision and recall, U-Net with EfficientNetB6 backbone performs well, suggesting the utility of a larger backbone with a bigger receptive field. However, ConvNeXt Small provides a trade-off between computational efficiency and performance, requiring fewer GFLOPs and parameters compared to UperNet with Swin Small while achieving similar performance.
@@ 5.1 Binary Segmentation
Architecture Backbone Params GFLOPs IoU F1 Score Precision Recall
U-Net ResNet50 32.52M 21.42 58$\pm$0.5 61.9$\pm$0.8 90.2$\pm$2.2 62.3$\pm$1.6
ResNext50 32M 21.81 58.2$\pm$0.2 62.1$\pm$0.3 88.2$\pm$3.5 63.6$\pm$1.7
SE_ResNet50 35.06M 20.83 58.9$\pm$0.7 62.9$\pm$0.7 88.8$\pm$1.6 64.4$\pm$1.4
EfficientNetB6 43.83M - 60.4$\pm$0.3 64.8$\pm$0.4 87.8$\pm$0.4 66.3$\pm$0.3
PAN ResNet50 24.26M 17.47 57.8$\pm$0.8 61.5$\pm$0.9 89.3$\pm$1.2 61.5$\pm$0.9
DeepLabV3+ 26.68M 18.44 56.8$\pm$0.7 60.6$\pm$0.7 89.4$\pm$1.3 61.8$\pm$1.1
Segformer mit-b0-ade 3.72M 3.52 57.6$\pm$0.5 61.3$\pm$0.6 82.6$\pm$2.9 69.2$\pm$2.1
UperNet ConvNexT-Small 128.29M 81.76 59.4$\pm$0.1 63.5$\pm$0.1 81.5$\pm$0.5 71.5$\pm$0.4
ConvNexT-Base 146.27M 121.99 59.7$\pm$0.3 63.8$\pm$0.2 81.1$\pm$0.7 72.2$\pm$0.2
swin small 81.15M 134.20 59.9$\pm$0.7 64.2$\pm$0.7 80.6$\pm$0.5 73.1$\pm$0.1
5.2 Binary Object Detection
@@
Architecture Backbone Params GFLOPs IoU0.1 IoU0.3 IoU0.5 mAP50 mAP50:95
RetinaNet ResNet50 18.87M 0.93 24.58$\pm$0.11 43.07$\pm$0.8 59.79$\pm$0.36 0.18$\pm$0.28 0.72$\pm$1.12
FasterRCNN 41.09M 24.7 36.79$\pm$1.07 46.95$\pm$0.66 61.29$\pm$0.35 5.20$\pm$1.00 19.12$\pm$3.41
FCOS 31.85M 25.81 34.79$\pm$0.99 48.51$\pm$0.59 62.66$\pm$0.43 9.67$\pm$1.47 30.46$\pm$3.11
SSD Lite MobileNet 3.71M 0.64 33.91$\pm$0.18 50.30$\pm$0.08 65.07$\pm$0.03 9.76$\pm$0.39 25.14$\pm$0.66
Table 6: Results for the binary segmentation task for U-Net with ResNet50 backbone, trained with and without NIR imagery. We find that NIR significantly improves performance.
Modality IoU F1 Score Precision Recall
RGB+NIR 58.00$\pm$0.50 61.9$\pm$0.80 90.20$\pm$2.20 62.30$\pm$1.60
RGB 56.60$\pm$0.44 60.50$\pm$0.35 87.00$\pm$1.40 62.54$\pm$0.13
Table 7: Results for object detection task for the FCOS Model with ResNet50 backbone evaluated over the test set for multiple input modality. We report the intersection over union (IoU) over thresholds $0.1,0.3,0.5$ and the mean average precision (mAP) for both IoU$=0.5$ and IoU$\in[0.5,0.95]$ thresholds.
Modality IoU0.1 IoU0.3 IoU0.5 mAP50 mAP50:95
RGB+NIR 34.79$\pm$0.99 48.51$\pm$0.59 62.66$\pm$0.43 9.67$\pm$1.47 30.46$\pm$3.11
RGB 32.39$\pm$2.88 46.80$\pm$2.07 61.23$\pm$1.58 5.7$\pm$3.65 20.00$\pm$10.40
Table 8: Performance comparison of U-Net with ResNet50 backbone for binary segmentation trained on active wells only versus all well types. Our results demonstrate that training merely on active wells is not sufficient to detect abandoned and suspended wells in practice – highlighting the importance of including all three types in the dataset.
Metric Train Set (Well Type Label Present) Test Set
IoU Active (I) 0.502
All (I+II+III) 0.576
F1 Score Active (I) 0.503
All (I+II+III) 0.614
Precision Active (I) 0.998
All (I+II+III) 0.913
Recall Active (I) 0.502
All (I+II+III) 0.614
@@
The observed gap in performance is likely due to the small size of the wells. It is well-known that single-stage CNN architectures (such as FCOS and SSD) often demonstrate better performance on small object detection than two-stage methods (such as Faster R-CNN), and this aligns with our observations. The exception here is the single-stage method RetinaNet, which, although it has comparable IoU scores, struggles to detect wells accurately.
SSD Lite stands out with the highest IoU0.5 score of $65.07\pm 0.03$ and IoU0.3 score of $50.3\pm 0.08$. Whereas all models are quite similar in terms of IoU0.1, the highest score by FasterRCNN is $36.79\pm 1.07$. Thus, SSD Lite and FCOS excel in localization, especially at higher IoU thresholds, while Faster R-CNN is adept at detecting objects with minimal overlap. All models demonstrate low performance in terms of mAP50, which assesses precision-recall trade-off and detection accuracy at an IoU threshold of 0.5. FCOS achieves the of $9.67\pm 1.47$ while SSD Lite achieves a score of $9.76\pm 0.39$. This may be due to these models not producing region proposals confidently enough, especially in instances with a large number of wells. Whereas over a broader evaluation with mAP50:95 which averages precision across IoU thresholds from 0.5 to 0.95. All models apart from RetinaNet provide much better results, with FCOS achieving the highest score of $30.46\pm 3.11$, indicating a decent performance in the identification of well instances. While IoU scores are decent with increasing thresholds, the mAP scores indicate that a more complex model may be required for this task.
5.3 Value Added from Dataset Breadth
In these experiments, we consider the value provided by our inclusion of (i) near-infrared (NIR) imagery in addition to RGB channels, and (ii) the inclusion of abandoned and suspended wells in addition to active wells.
5.3.1 Benefits of NIR imagery: (RGB v/s RGB+NIR)
We find that including the NIR band significantly enhances both object detection and segmentation performance compared to purely RGB data. In object detection (Table 7), RGB+NIR achieves higher Intersection over Union (IoU) scores across all thresholds (0.1, 0.3, and 0.5) and a considerable increase in mAP@50 (9.67$\pm$1.47 vs. 5.7$\pm$3.65) and mAP@50:95 (30.46$\pm$3.11 vs. 20$\pm$10.4). Similarly, in segmentation (Table 6), RGB+NIR shows superior performance in IoU (58$\pm$0.5 vs. 56.6$\pm$0.44), F1 Score (61.9$\pm$0.8 vs. 60.5$\pm$0.35), and Precision (90.2$\pm$2.2 vs. 87$\pm$1.4), while maintaining a slightly lower Recall (62.3$\pm$1.6 vs. 62.54$\pm$0.13). These improvements are due to the enhanced spectral information, which is particularly effective in identifying features such as ground depressions, often more visible in the NIR band.
5.3.2 Benefits of Using Multiple Well Types.
While much past work has focused on active wells, we show that training on such data is insufficient to detect abandoned and suspended wells (which are often significantly harder to e.g. degradation and vegetation growth over time). The results in Table 8 emphasize the necessity of including all well types (active, suspended, and abandoned) in the training dataset to ensure comprehensive detection across diverse scenarios. While training exclusively on active wells achieves high performance for individual well types, but it significantly under-performs in mixed-type contexts (IoU: 0.502), indicating poor generalization when all well types are present. Training exclusively on active wells significantly under-performs in mixed-type contexts (IoU: 0.502), indicating poor generalization when all well types are present. Conversely, training on all well types improves the model’s ability to handle real-world heterogeneity, as reflected by a higher IoU (0.576), F1 score (0.614), and recall (0.614) for the test set. The enhanced recall demonstrates the model’s capability to identify a broader range of wells, crucial for environmental monitoring, where missing even a single abandoned well could result in unaddressed methane emissions or groundwater contamination. Precision for active-only training (0.998) is higher than the model trained on all well types (0.913), but its inability to detect all wells in an image (due to low recall) limits its applicability. Therefore, incorporating all well types in training ensures balanced, reliable performance, allowing for accurate detection and classification in diverse and realistic contexts.
@@ 7 Conclusion
In this paper, we present the first large-scale benchmark dataset aimed at identifying oil and gas wells, with a focus on abandoned and suspended wells, which are a significant source of greenhouse gases and other pollutants.
We combine high-resolution imagery, an extensive database of well locations, and expert verification to create the Alberta Wells Dataset. We frame well identification both in terms of object detection and binary segmentation and evaluate the performance of a wide range of popular deep learning methods on these tasks. We find that the UperNet Model (with a ConvNexT backbone), in particular, represents the most promising baseline for the binary segmentation task, while for object detection, all models demonstrate more mixed results, with Single Stage Models (such as FCOS and SSD) providing a relatively promising baseline. These results show that the Alberta Wells Dataset represents both a challenging as well as a societally impactful set of tasks.
The value added by the dataset is twofold. First, most global fossil fuel-producing regions do not have databases of well locations comparable to that provided by AER. Alberta’s varied landscape provides an ideal setting for developing algorithms for the detection of wells, which can then be used directly in other locations or fine-tuned. Second, even in Alberta, the list developed by AER is not comprehensive, and many abandoned wells are believed to be missing; remote sensing algorithms can provide candidate locations for domain experts to examine so as to determine the true number and locations of abandoned wells in Alberta. Overall, we hope that our work will make it possible to (i) better estimate the total effects of well-related methane emissions (currently extremely uncertain [1]) and groundwater contamination, (ii) prioritize wells to be plugged and decontaminated. We believe that the scalability of remote sensing will make it an invaluable tool in mitigating the global environmental impact of abandoned oil and gas wells.
@@ 8 Acknowledgments and Disclosure of Funding
This research was made possible by the generous access to the Planet Lab subscription provided by Dr. Tim Elrick from McGill University’s Geography Department. We acknowledge the essential support from Planet Labs for their academic subscription and guidance on the public release of datasets. The Alberta Energy Regulator’s AER ST37 dataset was a crucial resource for our work.
The authors are also grateful to Mila and the Digital Research Alliance of Canada for the vital computing facilities, and to NVIDIA Corporation for their indispensable computational resources. Special thanks to Francis Pelletier, the lab’s Data Engineer, whose crucial feedback on our data collection pipeline and coding significantly improved our methods.
9 Impact Statement
This work presents the first large-scale machine learning benchmark for pinpointing onshore oil and gas wells, including abandoned, suspended, and active wells. Our dataset, the Alberta Wells Dataset, contains over 200,000 well instances and high-resolution satellite imagery, enabling advancements in the identification of oil and gas wells. By framing the problem as one of object detection and binary segmentation, we hope this dataset will help improve the scalability and accuracy of well detection techniques.
Abandoned wells, particularly those that leak methane and other pollutants, pose a major environmental challenge. By improving methods for detecting these wells, this work aims to reduce their environmental impact by identifying high-emission sites that can be prioritized for remediation. This can contribute to mitigating climate change by helping to reduce harmful emissions from these otherwise difficult-to-detect sources.
To enable wider responsible uses of the dataset, we will release it under a non-commercial usage license, making it available for academic and research purposes only. We do not foresee significant negative impacts arising from this work. Dual uses of the algorithms we outline here are limited, and the algorithms used are comparatively small-scale and have low energy and carbon footprint.
By improving detection techniques and providing access to valuable data, this work has the potential to support global efforts in identifying and addressing abandoned oil and gas wells, ultimately contributing to environmental sustainability and climate action.
Appendix A Data Licensing
The AWD Dataset is released under a Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) License (https://creativecommons.org/licenses/by-nc/4.0/). The satellite imagery for this project was acquired through Planet Labs’ [9] Education & Research license, which allows the use of the data in publications and the creation of derivative products related to those publications. However, the raw imagery cannot be shared publicly. To adhere to these guidelines, we provide the data in HDF5 format, with the satellite imagery pre-processed to produce a derived product represented as a numpy array from Raster Vector. This process removes all geographic metadata.
This data is for academic use only and should not be used commercially. Proper credit to the current authors, Planet Labs [9], and the Alberta Energy Regulator [31] is required when using this data.
Appendix B Dataset Information
@@
B.1 Dataset Structure
We provide training, validation, and testing sets, split using our proposed algorithm (as described in Section 3.2 of the main paper) to create a well-distributed dataset. The proposed method aims to create smaller regions of well concentration by clustering the centroids of patches. These regions are designed to be (a) mutually non-intersecting, (b) part of a larger geographic region by clustering the centroids of the initial clusters, and (c) containing a similar distribution of non-well patches within the same region.
This approach ensures that the training, validation, and test sets include representations from all geographic regions, providing a diverse and comprehensive evaluation. Thus, the dataset represents various geographical regions and offers a diverse benchmark for evaluation and testing.
@@ B.1 Dataset Structure
B.2 Dataset File Directory Structure
@@
    |---description: Alberta Wells Dataset: Pinpointing Oil and Gas Wells
@@
B.3 Structure of Dataset Directory
@@
    |---description: Alberta Wells Dataset: Pinpointing Oil and Gas Wells
@@
Dataset
Split
No of
Samples
No of Wells
in Split
Original HDF5
File Size (in Gb)
Compressed .tar.gz
File Size (in Gb)
Train 167436 194231 322 100
Validation 9463 8243 19 5.7
Test 11789 10973 24 7.1
Total 188688 213447 365 112.8
@@
B.4 Dataset Size & Distribution of Samples
Our dataset comprises over 94,000 patches of satellite imagery containing wells, with a total of 188,000 patches sourced from Planet Labs [9], covering more than 213,000 individual wells. Details about the distribution of the number of patches, wells present, and dataset split sizes are provided in Table 9, with the distribution of the number of wells per sample being described in Table 10. We also include an equal number of images that contain no wells in each dataset split. The distribution of wells per sample, along with the corresponding number of wells and the breakdown of well types, is illustrated in Figure LABEL:fig:figure_dist_act_sus_abd and detailed in Tables 11, 12, and 13 . The geographic distribution of wells in the dataset can be visualized in Figure 10.
Table 11: Test set statistics showing the distribution of image samples by the number of wells per image and the breakdown of well types (active, suspended, and abandoned).
No of Wells in a Image Sample (S) Test Set
Count (Image Samples) Distribution of Well Type in Samples
Total Wells Active Wells Suspended Wells Abandoned Wells
1 4128 4128 999 1325 1804
2 - 3 1242 2858 1042 844 972
4 - 5 328 1445 495 464 486
6 - 10 227 1662 604 555 503
11 - 15 38 471 144 181 146
16 - 25 19 352 56 184 112
26 - 35 2 57 0 56 1
10973 3340 3609 4024
Table 12: Validation Set statistics showing the distribution of image samples by the number of wells per image and the breakdown of well types (active, suspended, and abandoned).
No of Wells in a Image Sample (S) Validation Set
Count (Image Samples) Distribution of Well Type in Samples
Total Wells Active Wells Suspended Wells Abandoned Wells
1 3393 3393 743 1225 1425
2 - 3 979 2197 833 654 710
4 - 5 190 824 346 248 230
6 - 10 123 910 323 331 256
11 - 15 23 287 67 114 106
16 - 25 11 204 32 63 109
26 - 35 10 300 33 22 245
36 - 55 3 128 29 14 85
8243 2406 2671 3166
Table 13: Train Set statistics showing the distribution of image samples by the number of wells per image and the breakdown of well types (active, suspended, and abandoned).
No of Wells in a Image Sample (S) Train Set
Count (Image Samples) Distribution of Well Type in Samples
Total Wells Active Wells Suspended Wells Abandoned Wells
1 44299 44299 16715 13116 14468
2 - 3 25378 59190 33099 12706 13385
4 - 5 7899 34544 22456 6321 5767
6 - 10 4927 35333 21522 7796 6015
11 - 15 751 9345 4076 3136 2133
16 - 25 333 6381 1781 2544 2056
26 - 35 67 1994 376 791 827
36 - 55 45 1929 172 777 980
56 - 75 18 1129 86 345 698
76 - 125 1 87 11 63 13
83718 194231 100294 47595 46342
B.5 PlanetScope Satellite Imagery
For our experiments, we selected a 4-band (RGBN) satellite imagery product (ortho_analytic_4b_sr) from Planet Labs [9] as illustrated in Figure LABEL:fig:figure1a. This product uses Planet’s PSB.SD instrument, which features a telescope with a larger 47-megapixel sensor and is designed to be interoperable with Sentinel-2 imagery in several bands. The frequency of each band of image is described in Table 14. The instrument provides a frame size of 32.5 km x 19.6 km, an image capture capacity of 200 million km²/day, and an imagery bit depth of 12-bit, with a ground sample distance (nadir) ranging from 3.7 m to 4.2 m.
Figure 10: Density map of wells in the Alberta Wells Dataset.